find the solution y(t) to each of the following equations given the initial condition
3= (0) . - ܙ (b )
Answer #1

y= 24-2 using Seperable method, S (24-2) dy = fidt By2 2y = ttc y²ey=t+c y (6)=3 ie 9-6= Otc c=3. yery=t+3 y²2y+1=t+4 (4-1)²